OpenStudy (anonymous):

Determine the zeros of f(x) = x^3 – 12x^2 + 28x – 9.

OpenStudy (psymon):

Do we know about the rational zeros theorem at all?

OpenStudy (anonymous):

Yes

OpenStudy (psymon):

Okay, cool. So we need to use that to see what our possible zeros could be. If you know what that is then I'll just go through it. \[possible zeros = \frac{ \pm9, 3, 1 }{ \pm1 }\]. So know we need to choose one of these possible numbers and see if we can then do synthetic division and come up with a remainder of zero. Do we know how to do synthetic division?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do i test all of the factors now? If so I got, \[\pm1,\pm3,\pm9\]

OpenStudy (psymon):

Correct. We would need to test those out and hope we get a remainder of 0. I would try positive 9 first :P

OpenStudy (anonymous):

So my equation now is \[x ^{2}-3x+1\]?

OpenStudy (anonymous):

Now i find the discrimant and factor right?

OpenStudy (psymon):

Correct. Since you got a remainder of zero, this means (x-9) is a factor. And yes, you would then need to factor that portion.

OpenStudy (anonymous):

so what would my answer be?

OpenStudy (psymon):

Did you try to factor it?

OpenStudy (anonymous):

I didn't get real rumbers, what did i do wrong?

OpenStudy (psymon):

Well, I did it using the quadratic formula, so let me work it out for ya that way

OpenStudy (psymon):

\[\frac{ -(-3)\pm \sqrt{(-3)^{2}-4(1)(1)} }{ 2 }\] \[\frac{ 3\pm \sqrt{5} }{ 2 }\] So no, you should get real numbers.

OpenStudy (anonymous):

so what would be my solution?

OpenStudy (psymon):

\[(x-9) (x -\frac{ 3+\sqrt{5} }{ 2 }) (x-\frac{ 3 - \sqrt{5} }{ 2 })\] That would be how I'd personally put it. Unless you need to actually solve for x of course.

OpenStudy (anonymous):

So my zeroes are those answers?

OpenStudy (psymon):

Yes, those would be your zeros.

OpenStudy (anonymous):

thanks